# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v4

@@ 5 Conclusion
To provide readers with a more in-depth understanding of our LINO UniPS network architecture, in the following, we detail the architecture of our LINO UniPS, breaking it down into its seven key modules: (a) Wavelet DownSample Module, (b) Feature Extractor, (c) Light Registered Attention Module, (d) Fusion Block, (e) Wavelet UpSample Module, (f) Decoder and (g) the training loss. An overview of our network architecture is presented in Fig. 2. In the subsequent sections, we will provide a detailed account of the network’s structural components and the specific transformations of tensor shapes as data progresses through the model.
@@
Subsequently, both the downsample image representation $I_{f}^{d}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$ and the set of wavelet components $I_{f}^{w}$ (comprising $I_{f}^{ll},I_{f}^{lh},I_{f}^{hl},I_{f}^{hh}$, each in $\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$) are individually processed. First, each of these input components is partitioned into a sequence of patch-based tokens. These token sequences are then fed into our feature extraction backbone. This ViT backbone is trained during our training procedure to extract rich visual representations from these diverse inputs. In our specific implementation, we set the patch size to $P=8$. Consequently, for each input component with spatial dimensions $H/2\times W/2$, the resulting sequence length is $L=\frac{(H/2)\times(W/2)}{P^{2}}$ tokens. The feature embedding dimension is $D=384$. Following processing by this backbone, we respectively obtain shallow visual feature representations $F_{s,f}^{d}\in\mathbb{R}^{L\times D}$ from the downsample image stream (derived from $I_{f}^{d}$) and $F_{s,f}^{w}\in\mathbb{R}^{4\times L\times D}$ from the wavelet components stream (derived from $I_{f}^{w}$).
@@
Then $F_{s,f,r}^{d}$ and $F_{s,f,r}^{w}$ are fed into our Interleaved Attention Block Wang et al. (2025) to enhance inter-intra feature communication. Specifically, our Interleaved Attention Block contains four attention layers, which can be represented as: frame $\rightarrow$ light-axis $\rightarrow$ global $\rightarrow$ light-axis.
@@
Previous work has found that feature communication within the encoder is very important Ikehata (2022; 2023); Hardy et al. (2024), but their methods are often limited to patch-level local light-axis attention. Our alternating attention, however, breaks such limitations. On the one hand, it incorporates frame attention to enhance inter-image communication. On the other hand, we have added global attention, a more comprehensive global operation, allowing intra-image features to also be extended from the patch level to the image level. Upon processing by these four cascaded Interleaved Attention Blocks (the number chosen to maintain a manageable parameter count, although more blocks could potentially be employed), we obtain the deep feature representations denoted as $F_{d,f,r}^{d}\in\mathbb{R}^{L^{\prime}\times D}$ and $F_{d,f,r}^{w}\in\mathbb{R}^{4\times L^{\prime}\times D}$. It is worth noting that the attention operations inherent in these blocks do not alter the fundamental shapes of these tensor sequences.
@@
To ensure the Light Register Tokens $x_{\text{env}}$, $x_{\text{point}}$, $x_{\text{direction}}$ effectively capture global illumination, as direct supervision of the decoded light map is challenging, we introduce a light-aware feature alignment strategy during training Yu et al. (2025); Yao et al. (2025); Song et al. (2025). Since we train LINO UniPS on our own rendered synthetic dataset PS-Verse, for every scene within this training dataset, we have access to its corresponding lighting information from the rendering process. This includes: the HDRI environment map, the positions, the distance to camera and intensities of point lights, and the positions, the distance to camera, intensities and areas of direction lights. Mathematically, these are denoted as $L_{\text{env}}\in\mathbb{R}^{H\times W\times 3}$, $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 5}$, and $L_{\text{direction}}\in\mathbb{R}^{M_{2}\times 6}$, respectively. For $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 5}$, where $M_{1}$ is the number of point lights, its first three dimensions denote position, the fourth denotes the distance and the fifth denotes the intensity. For the directional light component $L_{\text{direction}}\in\mathbb{R}^{M_{2}\times 6}$, where $M_{2}$ denotes the number of directional lights, its first three dimensions specify position, the fourth denotes the distance, the fifth denotes the directional light size, and the sixth indicates intensity. Subsequently, we encode the lighting components $L_{\text{env}}$, $L_{\text{point}}$, and $L_{\text{direction}}$ by projecting them into a $D$-dimensional feature space, thereby obtaining their respective representations $\mathbf{l}^{h}_{\text{env}}\in\mathbb{R}^{D}$, $\mathbf{l}^{h}_{\text{point}}\in\mathbb{R}^{D}$, and $\mathbf{l}^{h}_{\text{direction}}\in\mathbb{R}^{D}$. Similarly, the light tokens $x_{\text{env}}$, $x_{\text{point}}$, and $x_{\text{direction}}$, after being processed by the cascaded Interleaved Attention Blocks, are projected into the same $D$-dimensional feature space. This yields their respective high-dimensional representations: $\mathbf{x}^{h}_{\text{env}}$, $\mathbf{x}^{h}_{\text{point}}$, and $\mathbf{x}^{h}_{\text{direction}}$, all in $\mathbb{R}^{D}$. Specifically, this projection is realized using three structurally similar two-layer Multi-Layer Perceptrons (MLPs). Within this common embedding space, we employ cosine similarity to supervise and align their respective feature distributions. Cosine similarity is chosen as it effectively measures the directional concordance between feature vectors, making it suitable for aligning representations of different lighting characteristics. This supervision translates into three distinct loss functions, denoted as $\mathcal{L}_{\text{env}}$, $\mathcal{L}_{\text{point}}$, and $\mathcal{L}_{\text{direction}}$. Their respective mathematical formulations are:
@@
Figure A1: Qualitative comparison on PSVerse Testdata. LINO UniPS demonstrates superior PBR estimation compared to all other methods.
@@
The subsequent discussion details the Fusion Block applied to features extracted during the initial stages of our Encoder. The overall feature fusion process is DPT-based Ranftl et al. (2021). For clarity, we will first illustrate this process using the features derived from the downsample image features $F_{d,f,r}^{d}$ as the primary example. Within each Interleaved Attention module of the encoder, features obtained from its four internal attention—frame, light-axis, global, and light-axis—are first concatenated along the feature dimension. This operation yields an aggregated feature set for each module, denoted as $F_{d,f}^{d,(i)}\in\mathbb{R}^{L\times 4D}$, where $i\in\{1,2,3,4\}$ represents the index of the $i$-th attention block. It is crucial to note that the three additional Light Register Tokens (introduced previously) do not participate in this specific feature aggregation (concatenation) process. Then, the aggregated features $F_{d,f}^{d,(i)}$ from four selected Interleaved Attention Blocks then undergo a series of projection and resizing operations. These operations transform them into a four-level feature pyramid, with hierarchical features $H_{1},H_{2},H_{3},$ and $H_{4}$. Their respective shapes are $F\times C\times H_{0}/2\times W_{0}/2$, $F\times 2C\times H_{0}/4\times W_{0}/4$, $F\times 4C\times H_{0}/8\times W_{0}/8$, and $F\times 4C\times H_{0}/16\times W_{0}/16$. In this context, $F$ represents the number of multi-light input images, $C$ is a base feature channel dimension (256), and $H_{0},W_{0}$ refer to the spatial resolution of the original, full-sized input images. These four levels of hierarchical features ($H_{1}$ through $H_{4}$) are subsequently progressively fused using residual convolutional blocks He et al. (2015), ultimately producing a fused feature representation $F_{\text{fused}}^{d}\in\mathbb{R}^{F\times\frac{H}{2}\times\frac{W}{2}\times C}$. A similar fusion process is applied to the features derived from the wavelet components path ($F_{d,f,r}^{w}$ ), yielding a corresponding fused representation, $F_{\text{fused}}^{w}\in\mathbb{R}^{4\times F\times\frac{H}{2}\times\frac{W}{2}\times C}$. It is noteworthy that this strategy of selecting four feature levels for hierarchical fusion can be adapted if more than four Interleaved Attention Blocks are employed in the encoder’s initial stages. For instance, if six such blocks are utilized, features from blocks indexed 1, 2, 4, and 6 might be selected to form the pyramid. Similarly, for an eight-block configuration, features from blocks 1, 3, 5, and 7 could be chosen as inputs to the hierarchical fusion pathway.
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}+\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}+\lambda_{4}\mathcal{L}_{\text{conf}}+\lambda_{5}\mathcal{L}_{\text{g}},$
@@
$\lambda_{1}=\frac{0.1}{(\mathcal{L}_{\text{env}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{2}=\frac{0.1}{(\mathcal{L}_{\text{point}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{3}=\frac{0.1}{(\mathcal{L}_{\text{direction}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{4}=1,\lambda_{5}=\frac{0.1}{(\mathcal{L}_{\text{g}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}$
@@
Our decision to set $\lambda_{4}$ to $1$ is because $\mathcal{L}_{\text{conf}}$ serves as the principal component in our overall loss function. The remaining auxiliary losses are then scaled using the adaptive weighting mechanism detailed in Eq. A1.1.7. This mechanism constrains their magnitudes to $0.1$ times that of the primary loss’s detached value, $(\mathcal{L}_{\text{conf}})_{\text{sg}}$, while still allowing their gradients to backpropagate fully. Such a strategy effectively positions these auxiliary losses to act as regularizers to the main learning objective, rather than allowing disparate loss magnitudes to vie for dominance and potentially destabilize training. This controlled weighting is crucial for ensuring stable and efficient training of LINO UniPS, mitigating issues such as excessively slow convergence or even training failure that can arise from an unbalanced multi-term loss function.
@@
Starting with our LINO UniPS model pre-trained for normal estimation, we finetune it on this complete PBR dataset for approximately 10 epochs until convergence. This process enables the model to simultaneously predict all four properties from the same input images. For the loss function, we augment the original loss $\mathcal{L}$ with a weighted MSE term for the material maps. The new loss is defined as:
$\mathcal{L_{\text{PBR}}}=\mathcal{L}+\lambda_{6}\mathcal{L}_{\text{a}}+\lambda_{7}\mathcal{L}_{\text{m}}+\lambda_{8}\mathcal{L}_{\text{r}}$
where $\mathcal{L}_{\text{a}},\mathcal{L}_{\text{m}},\mathcal{L}_{\text{r}}$ are the MSE between ground truth and predicted albedo, metallic, and roughness maps, respectively. The coefficients $\lambda_{6},\lambda_{7},\lambda_{8}$ are their corresponding loss weights, which are defined as:
$\lambda_{1}=\frac{1}{(\mathcal{L}_{\text{a}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{2}=\frac{1}{(\mathcal{L}_{\text{m}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{3}=\frac{1}{(\mathcal{L}_{\text{r}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}$
@@
Figure A3: The objects are displayed sequentially from left to right, representing Level 1 to Level 5. Across Levels 1 through 5, there is a progressive rise in geometric complexity. Specifically, Level 5 features exceptionally complex surface geometry due to the utilization of normal maps in its rendering process.
@@
To rigorously evaluate and enhance the capability of our LINO UniPS method for reconstructing surface normals of objects that feature high-frequency geometric details on complex surfaces, we curated a dedicated set of objects exhibiting diverse geometric complexities. These objects were subsequently graded by difficulty into five distinct levels, designated Level 1 to Level 5. Specifically, Levels 1–4 are classified following Dora Chen et al. (2024) criterion, based on the number of salient edges $N_{\Gamma}$. Level 5, in contrast, is distinguished by the use of normal maps in its rendering. The specific criteria for this classification are as follows:
@@
  • Level 5 : With Normal Map.
Fig. A3 shows representative cases from the different defined levels. PS-Verse comprises 100,000 scenes. For each of these scenes, two distinct renderings are typically generated: one that utilizes normal map to incorporate fine geometric details, and another rendered without this normal map. Levels 1 4 consist exclusively of scenes rendered without normal map, with each of these four levels containing 25,000 scenes. Level 5 is composed entirely of the 100,000 scenes rendered with normal map enhancement.
Figure A4: Effect of normal mapping on rendered surface detail. The top two rows display renderings without normal maps, while the bottom two rows showcase the same scenes rendered with normal map. It is evident that employing normal maps (bottom rows) results in significantly more high-frequency surface normal detail compared to renderings without (top rows).
A1.3.2 The Use of Normal Map
To enhance PS normal reconstruction for objects characterized by intricate, high-frequency details, training data rich in such geometric features is essential. However, 3D models genuinely possessing fine-grained geometric intricacies are often scarce and prohibitively expensive, which impedes the creation of diverse, large-scale, high-fidelity datasets. To overcome this limitation within the Universal PS framework, our work pioneers the integration of normal mapping directly into the dataset generation process. Normal mapping, a 3D computer graphics technique, imbues low-polygon models with the visual appearance of high-frequency geometric details by applying a specialized texture, known as a normal map, which encodes fine-scale perturbations of the surface normals. During rendering, these stored normal variations are then utilized to simulate intricate surface details without actually increasing the underlying geometric complexity or polygon count of the model.
A visual comparison of renderings with and without the use of normal maps is presented in Fig. A4. It is clearly evident from the figure that employing normal maps during the rendering process yields a significantly higher level of detail in the resulting surface normals.
@@
During rendering, we generate ten distinct lighting configurations by combining several base lighting components (conceptually denoted here as (a), (b), (c), and (d)). These specific configurations are as follows: (1) Component (a), (2) Component (b), (3) Component (b), (4) Components (a) + (b), (5) Components (a) + (c), (6) Components (b) + (c), (7) Components (a) + (b) + (c), (8) Components (a) + (d), (9) Components (b) + (d), (10) Components (a) + (b) + (d). The lighting setup includes: directional light, point light and uniform background lighting, which is introduced to better simulate realistic global illumination. Every scene in PS-Verse is rendered as 20 images, each employing a lighting setup randomly chosen from our ten predefined lighting configurations. An example of such an image set for a single scene is illustrated in Fig. A2.
@@
Here, we primarily compare several training datasets: CyclesPS-Train Ikehata (2018), PS-Wild Ikehata (2022), PS-Mix Ikehata (2023), PS-Uni MS-PS Hardy et al. (2024), and our PS-Verse. For a quantitative comparison of these datasets, please refer to Tab. 1. We now present illustrative qualitative comparisons in Fig. A5. Our comprehensive evaluation, encompassing both qualitative and quantitative aspects, leads us to conclude that PS-Verse is the premier training dataset in terms of quality for the Universal PS task.
@@
Figure A7: Post-PCA visualization of features extracted by different method encoders for the CowPNG from the DiLiGenT Shi et al. (2018). Metrics displayed to the right of each row is (CSIM/SSIM), higher values indicate higher feature similarity.

### v5

@@ 5 Conclusion
Acknowledgements.
This work was partially supported by a grant from the NSFC/RGC Collaborative Research Scheme sponsored by the Research Grants Council of the Hong Kong Special Administrative Region, China and National Natural Science Foundation of China (Project No. CRS_HKUST605/25).
@@ Acknowledgements.
Figure A1: An overview of our network architecture, illustrating the corresponding tensor shape transformations. Black solid arrows denote the forward pass of the downsample branch, while purple dashed arrows denote the forward pass of the wavelet branch.
To provide readers with a more in-depth understanding of our LINO UniPS network architecture, in the following, we detail the architecture of our LINO UniPS, breaking it down into several key modules: (a) Wavelet DownSample Module, (b) Feature Extractor, (c) Light Registered Attention Module, (d) Fusion Block, (e) Wavelet UpSample Module, (f) Decoder and the training loss. An overview of our network architecture is presented in Fig. A1. In the subsequent sections, we will provide a detailed account of the network’s structural components and the specific transformations of tensor shapes as data progresses through the model.
@@
Subsequently, both the downsample image representation $I_{f}^{d}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$ and the set of wavelet components $I_{f}^{w}$ (comprising $I_{f}^{ll},I_{f}^{lh},I_{f}^{hl},I_{f}^{hh}$, each in $\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 3}$) are individually processed. First, each of these input components is partitioned into a sequence of patch-based tokens. These token sequences are then fed into our Feature Extractor backbone. This backbone is trained during our training procedure to extract rich visual representations from these diverse inputs. In our specific implementation, we set the patch size to $P=8$. Consequently, for each input component with spatial dimensions $H/2\times W/2$, the resulting sequence length is $L=\frac{(H/2)\times(W/2)}{P^{2}}$ tokens. The feature embedding dimension is $D=384$. Following processing by this backbone, we respectively obtain shallow visual feature representations $F_{s,f}^{d}\in\mathbb{R}^{L\times D}$ from the downsample image stream (derived from $I_{f}^{d}$) and $F_{s,f}^{w}\in\mathbb{R}^{4\times L\times D}$ from the wavelet components stream (derived from $I_{f}^{w}$).
@@
Then $F_{s,f,r}^{d}$ and $F_{s,f,r}^{w}$ are fed into our Interleaved Attention Block Wang et al. (2025) to enhance inter-intra feature communication. Specifically, our Interleaved Attention Block contains four attention layers, which can be represented as: Frame $\rightarrow$ Light $\rightarrow$ Global $\rightarrow$ Light.
@@
Previous work has found that feature communication within the encoder is very important Ikehata (2022; 2023); Hardy et al. (2024), but their methods are often limited to patch-level local light-axis attention. Our Interleaved Attention Block, however, breaks such limitations. On the one hand, it incorporates Frame attention to enhance intra-image communication. On the other hand, we have added Global attention, a more comprehensive global operation, allowing inter-image features to also be extended from the patch level to the image level. Upon processing by these four cascaded Interleaved Attention Blocks (the number chosen to maintain a manageable parameter count, although more blocks could potentially be employed), we obtain the deep feature representations denoted as $F_{d,f,r}^{d}\in\mathbb{R}^{L^{\prime}\times D}$ and $F_{d,f,r}^{w}\in\mathbb{R}^{4\times L^{\prime}\times D}$. It is worth noting that the attention operations inherent in these blocks do not alter the fundamental shapes of these tensor sequences.
@@
To ensure the Light Register Tokens $x_{\text{env}}$, $x_{\text{point}}$, $x_{\text{direction}}$ effectively capture global illumination, as direct supervision of the decoded light map is challenging, we introduce a light-aware feature alignment strategy during training Yu et al. (2025); Yao et al. (2025); Song et al. (2025). Since we train LINO UniPS on our own rendered synthetic dataset PS-Verse, for every scene within this training dataset, we have access to its corresponding lighting information from the rendering process. This includes: the HDRI environment map, the positions, the distance to camera and intensities of point lights, and the positions, the distance to camera, intensities and areas of directional lights. Mathematically, these are denoted as $L_{\text{env}}\in\mathbb{R}^{H\times W\times 3}$, $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 5}$, and $L_{\text{direction}}\in\mathbb{R}^{M_{2}\times 6}$, respectively. For $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 5}$, where $M_{1}$ is the number of point lights, its first three dimensions denote position, the fourth denotes the distance and the fifth denotes the intensity. For the directional light component $L_{\text{direction}}\in\mathbb{R}^{M_{2}\times 6}$, where $M_{2}$ denotes the number of directional lights, its first three dimensions specify position, the fourth denotes the distance, the fifth denotes the directional light size, and the sixth indicates intensity. Subsequently, we encode the lighting components $L_{\text{env}}$, $L_{\text{point}}$, and $L_{\text{direction}}$ by projecting them into a $D$-dimensional feature space, thereby obtaining their respective representations $\mathbf{l}^{h}_{\text{env}}\in\mathbb{R}^{D}$, $\mathbf{l}^{h}_{\text{point}}\in\mathbb{R}^{D}$, and $\mathbf{l}^{h}_{\text{direction}}\in\mathbb{R}^{D}$. Similarly, the light tokens $x_{\text{env}}$, $x_{\text{point}}$, and $x_{\text{direction}}$, after being processed by the cascaded Interleaved Attention Blocks, are projected into the same $D$-dimensional feature space. This yields their respective high-dimensional representations: $\mathbf{x}^{h}_{\text{env}}$, $\mathbf{x}^{h}_{\text{point}}$, and $\mathbf{x}^{h}_{\text{direction}}$, all in $\mathbb{R}^{D}$. Specifically, this projection is realized using three structurally similar two-layer Multi-Layer Perceptrons (MLPs). Within this common embedding space, we employ cosine similarity to supervise and align their respective feature distributions. Cosine similarity is chosen as it effectively measures the directional concordance between feature vectors, making it suitable for aligning representations of different lighting characteristics. This supervision translates into three distinct loss functions, denoted as $\mathcal{L}_{\text{env}}$, $\mathcal{L}_{\text{point}}$, and $\mathcal{L}_{\text{direction}}$. Their respective mathematical formulations are:
@@
Figure A2: Qualitative comparison on PS-Verse Testdata. LINO UniPS demonstrates superior PBR estimation compared to all other methods.
@@
The subsequent discussion details the Fusion Block applied to features extracted during the initial stages of our Encoder. The overall feature fusion process is DPT-based Ranftl et al. (2021). For clarity, we will first illustrate this process using the features derived from the downsample image features $F_{d,f,r}^{d}$ as the primary example. Within each Interleaved Attention Block of the encoder, features obtained from its four internal attention—Frame, Light Axis, Global, and Light Axis—are first concatenated along the feature dimension. This operation yields an aggregated feature set for each block, denoted as $F_{d,f}^{d,(i)}\in\mathbb{R}^{L\times 4D}$, where $i\in\{1,2,3,4\}$ represents the index of the $i$-th attention block. It is crucial to note that the three additional Light Register Tokens (introduced previously) do not participate in this specific feature aggregation (concatenation) process. Therefore, the feature dimension of $F_{d,f}^{d,(i)}$ is $4D$, not $4D^{\prime}$.
Then, to effectively fuse features from different depths within the encoder, our Fusion Block employs a top-down, multi-scale fusion strategy. This process begins by selecting the aggregated output features, $F_{d,f}^{d,(i)}$, from four different stages of the Interleaved Attention Blocks. These features are initially 1D token sequences. First, these token sequences are reshaped from their 1D sequence format back into 2D spatial feature maps. Next, these four distinct spatial maps (which come from different depths of the Interleaved Attention Blocks) undergo a series of projection and downsampling operations to transform them into a standardized, four-level hierarchical feature pyramid, denoted $H_{1},H_{2},H_{3},$ and $H_{4}$. These operations typically consist of 1x1 convolutions to project the features to their target channel dimensions ($C,2C,4C,4C$) and downsampling (2x2 convolutions with a stride of 2) to match the target resolutions.This pyramid captures multi-scale information, ranging from high-resolution features at $H_{1}$ (shape $F\times C\times H/2\times W/2$) to low-resolution/high-semantic features at $H_{4}$ (shape $F\times 4C\times H/16\times W/16$), where $F$ is the number of multi-light images, $C$ is 256, and $H,W$ refer to the spatial resolution of the original, full-sized input images. We then employ a progressive, top-down fusion path. This fusion is implemented using residual convolutional blocks He et al. (2015) and upsampling operations (2x2 transposed convolution with a stride of 2). Specifically, the deepest feature $H_{4}$ is first upsampled and then fused with $H_{3}$ via a residual block; this result is then upsampled and fused with $H_{2}$, and so on.The final output of this progressive fusion is a single, information-rich fused feature map, $F_{\text{fused}}^{d}$, with a shape of $\mathbb{R}^{F\times\frac{H}{2}\times\frac{W}{2}\times C}$, which is then passed to the Wavelet UpSample module. A similar fusion process is applied to the features derived from the wavelet components path ($F_{d,f,r}^{w}$ ), yielding a corresponding fused representation, $F_{\text{fused}}^{w}\in\mathbb{R}^{4\times F\times\frac{H}{2}\times\frac{W}{2}\times C}$. It is noteworthy that this strategy of selecting four feature levels for hierarchical fusion can be adapted if more than four Interleaved Attention Blocks are employed in the encoder’s initial stages. For instance, if six such blocks are utilized, features from blocks indexed 1, 2, 4, and 6 might be selected to form the pyramid. Similarly, for an eight-block configuration, features from blocks 1, 3, 5, and 7 could be chosen as inputs to the hierarchical fusion pathway.
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}+\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$ (A2)
@@
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}+\lambda_{4}\mathcal{L}_{\text{conf}}+\lambda_{5}\mathcal{L}_{\text{g}}$ (A3)
@@
$\begin{gathered}\lambda_{1}=\frac{0.1}{(\mathcal{L}_{\text{env}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{2}=\frac{0.1}{(\mathcal{L}_{\text{point}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\\ \lambda_{3}=\frac{0.1}{(\mathcal{L}_{\text{direction}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{4}=1,\lambda_{5}=\frac{0.1}{(\mathcal{L}_{\text{g}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}\end{gathered}$ (A4)
@@
Our decision to set $\lambda_{4}$ to $1$ is because $\mathcal{L}_{\text{conf}}$ serves as the principal component in our overall loss function. The remaining auxiliary losses are then scaled using the adaptive weighting mechanism detailed in Eq. A4. This mechanism constrains their magnitudes to $0.1$ times that of the primary loss’s detached value, $(\mathcal{L}_{\text{conf}})_{\text{sg}}$, while still allowing their gradients to backpropagate fully. Such a strategy effectively positions these auxiliary losses to act as regularizers to the main learning objective, rather than allowing disparate loss magnitudes to vie for dominance and potentially destabilize training. This controlled weighting is crucial for ensuring stable and efficient training of LINO UniPS, mitigating issues such as excessively slow convergence or even training failure that can arise from an unbalanced multi-term loss function.
@@
[…]
$\lambda_{6}=\frac{1}{(\mathcal{L}_{\text{a}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{7}=\frac{1}{(\mathcal{L}_{\text{m}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{8}=\frac{1}{(\mathcal{L}_{\text{r}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}$ (A5)
@@
Figure A4: The objects are displayed sequentially from left to right, representing Level 1 to Level 5. Across Levels 1 through 5, there is a progressive rise in geometric complexity. Specifically, Level 5 features exceptionally complex surface geometry due to the utilization of normal mapping in its rendering process.
@@
To rigorously evaluate and enhance the capability of our LINO UniPS method for reconstructing surface normals of objects that feature high-frequency geometric details on complex surfaces, we curated a dedicated set of objects exhibiting diverse geometric complexities. These objects were subsequently graded by difficulty into five distinct levels, designated Level 1 to Level 5. Specifically, Levels 1–4 are classified following Dora Chen et al. (2024) criterion, based on the number of salient edges $N_{\Gamma}$. Level 5, in contrast, is distinguished by the use of normal mapping in its rendering. The specific criteria for this classification are as follows:
@@
  • Level 5 : With Normal Mapping
Fig. A4 shows representative cases from the different defined levels. PS-Verse comprises 100,000 scenes. For each of these scenes, two distinct renderings are typically generated: one that utilizes normal mapping to incorporate fine geometric details, and another rendered without this normal mapping. Levels 1-4 consist exclusively of scenes rendered without normal mapping, with each of these four levels containing 25,000 scenes. Level 5 is composed entirely of the 100,000 scenes rendered with normal mapping enhancement.
Figure A5: Effect of normal mapping on rendered surface detail. The top two rows display renderings without normal mapping, while the bottom two rows showcase the same scenes rendered with normal mapping. It is evident that employing normal mapping (bottom rows) results in significantly more high-frequency surface normal detail compared to renderings without (top rows).
A1.3.2 The Use of Normal Mapping
To enhance PS normal reconstruction for objects characterized by intricate, high-frequency details, training data rich in such geometric features is essential. However, 3D models genuinely possessing fine-grained geometric intricacies are often scarce and prohibitively expensive, which impedes the creation of diverse, large-scale, high-fidelity datasets. To overcome this limitation within the Universal PS framework, our work pioneers the integration of normal mapping directly into the dataset generation process. Normal mapping, a 3D computer graphics technique, imbues low-polygon models with the visual appearance of high-frequency geometric details by applying a specialized texture, which encodes fine-scale perturbations of the surface normals. During rendering, these stored normal variations are then utilized to simulate intricate surface details without actually increasing the underlying geometric complexity or polygon count of the model.
A visual comparison of renderings with and without the use of normal mapping is presented in Fig. A5. It is clearly evident from the figure that employing normal mapping during the rendering process yields a significantly higher level of detail in the resulting surface normals.
@@
During rendering, we generate ten distinct lighting configurations by combining several base lighting components (conceptually denoted here as (a), (b), (c), and (d)). These specific configurations are as follows: (1) Component (a), (2) Component (b), (3) Component (c), (4) Components (a) + (b), (5) Components (a) + (c), (6) Components (b) + (c), (7) Components (a) + (b) + (c), (8) Components (a) + (d), (9) Components (b) + (d), (10) Components (a) + (b) + (d). The lighting setup includes: directional light, point light and uniform background lighting, which is introduced to better simulate realistic global illumination. Every scene in PS-Verse is rendered as 20 images, each employing a lighting setup randomly chosen from our ten predefined lighting configurations. An example of such an image set for a single scene is illustrated in Fig. A3.
@@
Here, we primarily compare several training datasets: CyclesPS-Train Ikehata (2018), PS-Wild Ikehata (2022), PS-Mix Ikehata (2023), PS-Uni MS-PS Hardy et al. (2024), and our PS-Verse. For a quantitative comparison of these datasets, please refer to Tab. 1. We now present illustrative qualitative comparisons in Fig. A6. Our comprehensive evaluation, encompassing both qualitative and quantitative aspects, leads us to conclude that PS-Verse is the premier training dataset in terms of quality for the Universal PS task. This emphasis on large-scale, high-quality data curation is also consistent with recent trends in dense correspondence and stereo matching Guo et al. (2025b; a).
@@
Figure A8: Post-PCA visualization of features extracted by different method encoders for the CowPNG from the DiLiGenT Shi et al. (2018). Metrics displayed to the right of each row are (CSIM/SSIM), higher values indicate higher feature similarity.